OpenStudy (anonymous):

1. 2x + 3y = -3 Which of the following states the x-intercept and the y-intercept of the equation above? a. x-intercept = (-1 1/2, 0) y-intercept = (0, -1) b. x-intercept = (0, -2/3) y-intercept = (-1, 0) c. x-intercept = (5, 0) y-intercept = (0, 6) d. x-intercept = (0, -1 1/2) y-intercept = (-1, 0)

OpenStudy (hoblos):

to find the x-intercept.. replace y=0 and solve for x to find the y-intercept.. replace x=0 and solve for y for y=0 2x=-3 x=-3/2 = -(1 1/2) so x intercept is (-1 1/2, 0) for x=0 3y=-3 y=-1 so y intercept is (0, -1)